The math contest called the Fermat is out of 150 points. The mean score this year is 91 points with a standard deviation of 12. Students in the top 25% receive a medal. Assuming the scores are normally distributed, what mark is required to receive a medal? *




1
Expert's answer
2021-07-26T15:42:35-0400

"P(Z<z)=0.75."

"z=0.67."

"\frac{x-91}{12}=0.67."

"x=91+0.67*12=99."

"P(X>99)=1-P(X<99)=0.25."